# City-on-Web: Real-time Neural Rendering of Large-scale Scenes on the Web

### v1

@@ Abstract
NeRF has significantly advanced 3D scene reconstruction, capturing intricate details across various environments. Existing methods have successfully leveraged radiance field baking to facilitate real-time rendering of small scenes. However, when applied to large-scale scenes, these techniques encounter significant challenges, struggling to provide a seamless real-time experience due to limited resources in computation, memory, and bandwidth. In this paper, we propose City-on-Web, which represents the whole scene by partitioning it into manageable blocks, each with its own Level-of-Detail, ensuring high fidelity, efficient memory management and fast rendering. Meanwhile, we carefully design the training and inference process such that the final rendering result on web is consistent with training. Thanks to our novel representation and carefully designed training/inference process, we are the first to achieve real-time rendering of large-scale scenes in resource-constrained environments. Extensive experimental results demonstrate that our method facilitates real-time rendering of large-scale scenes on a web platform, achieving 32FPS at 1080P resolution with an RTX 3060 GPU, while simultaneously achieving a quality that closely rivals that of state-of-the-art methods. Project page: https://ustc3dv.github.io/City-on-Web/.
@@ 1 Introduction
NeRF has significantly advanced the field of scene reconstruction, showing an unparalleled ability to capture complex details across diverse environments. Existing works have demonstrated its ability to render small scenes with exceptional quality and performance in real-time [3, 46, 47, 36, 21, 29, 6, 40, 11, 43, 48, 24]. NeRF has also been successfully applied to the rendering of large scenes in offline settings, achieving exceptional visual fidelity and generating intricately detailed results [34, 14, 43, 45, 37].
Despite these successes, real-time neural rendering of large scenes is profoundly challenging due to inherent computational power, memory, and bandwidth limitations across various devices. The challenges mainly include the following aspects. Firstly, traditional NeRF and its variants are resource-intensive, requiring substantial computational power that exceeds what is typically available in such constrained environments. Secondly, the video memory capacity on client devices is frequently limited, imposing significant restrictions on the capability to process and render substantial assets in real-time simultaneously. The substantial resource becomes a critical issue in the real-time rendering of large scenes, necessitating the quick loading and processing of extensive data sets. Lastly, the dependency on data retrieval from remote servers introduces latency, particularly under network bandwidth limitations, further complicating the real-time rendering process. These hurdles collectively form a significant barrier to delivering an uninterrupted and instantaneous visual experience for large-scale scenes.
To address these challenges in real-time rendering of large-scale scenes, we propose our proposed method, City-on-Web. Drawing inspiration from traditional graphics techniques for rendering large-scale scenes [10, 38, 7, 9, 15, 8, 23], we partition the scene into manageable blocks and represent the scene with varying Levels-of-Detail (LOD). We utilize radiance field baking techniques [16, 30], which precompute and store rendering primitives into 3D atlas textures organized in a sparse grid within each block for real-time rendering. However, due to the unavoidable texture resource limitations of shaders, we cannot load all the atlas textures into a single shader. Hence, we represent the scene as a hierarchy of segmented blocks, each rendered by a dedicated shader during rendering.
@@
2 Related Work
Figure 1: Pipeline of training. (a) We divide the entire scene into different blocks without overlap according to the ground plane. (b) For block $k$ that the ray $\boldsymbol{r}(t)$ passes through, the corresponding shader renders the color $\boldsymbol{c}^{k}$ and opacity $\alpha^{k}$ of each block. We depth sort the blocks that ray traversed, and then render the final result through alpha blending that maintains 3D consistency.
Large-scale Scene Reconstruction.
For radiance field reconstruction of large-scale scenes, a key issue lies in enhancing the model’s representational capacity to adequately capture and render extensive scenes. Block-NeRF [34] and Mega-NeRF [37] address this by adopting a divide-and-conquer strategy, segmenting expansive scenes into smaller blocks, and applying localized NeRF processing to each. This approach significantly improves both the reconstruction quality and the model’s scalability to larger scenes. Switch-NeRF [50] employs a gating network to dispatch 3D points to different NeRF sub-networks. Grid-NeRF [45] utilizes a compact multiresolution feature plane and combines the strengths of smoothness from vanilla NeRF with the local detail capturing ability of feature grid-based methods [26, 5, 31], efficiently reconstructing large scenes with fine details. NeRF++ [48] enhances the reconstruction of unbounded scenes through its innovative multi-spherical representation. On the other hand, Mip-NeRF 360 [1] introduces a scene contraction function to effectively represent scenes that extend to infinity, addressing the challenge of vast spatial extents. F2-NeRF [41] takes this a step further by implementing a warping function for local spaces, ensuring a balance of computational resources and training data across different parts of the scene.
Real-time Rendering.
Early works mainly focus on the real-time rendering of a simple single object. NSVF [21] improves NeRF by introducing a more efficient sparse voxel field, significantly accelerating rendering speed while maintaining high-quality output. KiloNeRF [29] utilizes thousands of small MLPs, each responsible for a tiny scene region, significantly reducing network evaluation time. In contrast, SNeRG [16] leverages pre-computed sparse grids, allowing for direct retrieval of radiance field information without needing network evaluation. Termi-NeRF [28] terminates ray marching in less impactful scene regions, slashing computation time. DONeRF [27] focuses on one sample using a depth oracle network, speeding up rendering while preserving scene quality. Recently, there have been developments that enable real-time rendering of neural radiance fields in small scenes. MERF [30] improves upon SNeRG by utilizing a voxel and triplane hybrid representation to reduce memory usage. MobileNeRF [6] introduces the polygon rasterization rendering pipeline, running NeRF-based novel view synthesis in real-time on mobile devices. BakedSDF [46] bakes volumetric representation into meshes and utilizes spherical harmonics for representing view-dependent color, while NeRF2Mesh [36] iteratively refine both the geometry and appearance of the mesh.
Level of Detail.
Substantial works are devoted to integrating LOD methods into the fabric of traditional computer graphics [7, 8, 9, 15, 20, 23, 17, 22], aiming to streamline rendering processes, reduce memory footprint, bolster interactive responsiveness. Recently, some works begin to apply LOD to the neural implicit reconstruction. NGLoD [32] represents LOD through a sparse voxel octree, where each level of the octree corresponds to a different LOD, allowing for a finer discretization of the surface and more detailed reconstruction as the tree depth increases. Takikawa et al. [33] efficiently encode 3D signals into a compact, hierarchical representation using vector-quantized auto decoder method. BungeeNeRF [43] employs a hierarchical network structure, where the base network focuses on learning a coarse representation of the scene, and subsequent residual blocks are tasked with progressively refining this representation. TrimipRF [18] and LoD-Neus [51] leverage multi-scale triplane and voxel representations to capture scene details at different scales, effectively implementing anti-aliasing to enhance the rendering and reconstruction quality.
3 Background and Motivation
Our exploration begins with an in-depth analysis of two influential works, SNeRG [16] and MERF [30], which have both set benchmarks for real-time rendering of radiance field. SNeRG precomputes and stores a Neural Radiance Fields model in a sparse 3D voxel grid. Each active voxel in SNeRG contains several attributes: density, diffuse color, and specular feature vector that captures view-dependent effects. Additionally, an indirection grid is used to enhance rendering by either indicating empty macroblocks or pointing to detailed texels in a 3D texture atlas. This representation allows real-time rendering on standard laptop GPUs.
The indirection grid assists in raymarching through the sparse 3D grid by passing empty regions and selectively accessing non-zero densities $\sigma_{i}$, diffuse colors $\boldsymbol{c}_{i}$, and feature vectors $\boldsymbol{f}_{i}$ during rendering. Integrating along each ray $\boldsymbol{r}(t)=\boldsymbol{o}+t\boldsymbol{d}$, we compute the sum of the weights, which can be considered as the pixel’s opacity:
The color $\boldsymbol{C}_{d}(\boldsymbol{r})$ and specular feature $\boldsymbol{F}_{s}(\boldsymbol{r})$ along the ray are accumulated using the same weights to compute the final diffuse color and specular feature of ray:
The step size during ray marching $\delta_{i}$ is equal to the voxel width for an occupied voxel. Subsequently, the accumulated diffuse color and specular feature vector, along with the positional encoding $PE(\cdot)$ of the ray’s view direction, are concatenated to pass through a lightweight deferred MLP $\Phi$ to produce a view-dependent residual color:
While SNeRG achieves impressive real-time rendering results, its voxel representation demands substantial memory, which poses limitations for further applications. MERF presents a significant reduction in memory requirements in comparison to extant radiance field methods like SNeRG. By leveraging hybrid low-resolution sparse grid and 2D high-resolution triplanes, MERF optimizes the balance between performance and memory efficiency. Moreover, it incorporates two pivotal strategies to bridge the gap between training and rendering performance. First, MERF simulates finite grid approach during training, querying MLPs at virtual grid corners and applying interpolation to mimic the rendering process closely. Second, MERF simulates quantization during training and employs the straight-through estimator [2], allowing for the simulation of the quantization process while maintaining differentiability, enabling the model to learn and optimize with quantized values without introducing non-differentiable steps during the backward pass, ensuring a smooth training process.
These innovative methods for scene reconstruction offer promising results, but their direct applicability to large scenes remains challenges. MERF’s hybrid voxel-triplane representation, despite being memory-efficient, cannot capture large scenes with intricate details due to its fixed resolution constraint. In our efforts to reconstruct large scenes with high fidelity, dividing them into smaller blocks is a practical solution. This method helps in making sure the reconstruction is detailed and accurate.
However, this approach of dividing scenes means we will end up with more assets to deal during rendering. Web browsers have limits on how much memory they can use, which can make it hard to show these many pieces of large and detailed models. Another challenge is the data transmission during web rendering. It often requires pulling data from servers, and this can be slow due to network delays. As a result, users might experience long wait times when trying to load all the detailed pieces of a large scene at once for rendering. Drawing inspiration from traditional mesh dynamic resource loading and LOD [39], we generate LOD from our segmented reconstruction results which helps in minimizing the load of distant resources. Additionally, by employing a dynamic loading strategy for blocks, we significantly reduce VRAM usage and decrease the wait time for resource transmission.
4 Method
In this section, we present a method for representing and rendering large scenes on the web. Our approach uses hierarchical spatial partitioning and LOD to manage large-scale scenes dynamically (Sec. 4.1). We align the training and rendering stages to ensure consistency (Sec. 4.2), employing multiple shaders and alpha blending for seamless integration of scene blocks. Additionally, the framework includes optimization strategies (Sec. 4.3) and a process for generating LODs (Sec. 4.4) and baking the model (Sec. 4.5) for real-time rendering.
4.1 Large-scale Radiance Field
In the realm of scene reconstruction and rendering, NeRF has made significant strides, achieving compelling results. However, it faces inherent challenges when tasked with representing large scenes on the web. Using a single model to represent such vast scenes proves challenging due to its limited expressiveness, particularly in achieving a detailed and accurate reconstruction. Representing scenes with multiple models in a single resolution increases the overhead during rendering, leading to the loading of numerous resources, which is not conducive to efficient rendering.
To efficiently represent large scenes captured using the fly-through method, we employ hierarchical spatial partitioning combined with LOD to represent the scene. Specifically, we uniformly partition the area into varying blocks within our region of interest on the $xy$ plane (i.e., the ground plane). Each set of partitioned blocks corresponds to a unique LOD level, allowing for dynamic and efficient representation. Within each block, we use a low resolution voxel with a high resolution triplane that stores density, diffuse color, and specular color, to represent the radiance field for web rendering. Additionally, for blocks along the periphery, we utilize the scene contraction function from MERF [30] to account for the data on the boundaries. For internal blocks, we simply adopt settings from the bounded scene.
During the training stage, our scene representation aligns with web rendering. We consistently utilize spatial partitioning to structure the scene into distinct blocks without overlap. However, we train the scene only with the finest level of LOD. Within block $k$ , the following trainable components are introduced: (1) $f^{k}$: an attribute query function which adopts a hash encoding and a mlp decoder that outputs attributes of points such as densities, diffuse color and specular feature (2) $\Phi^{k}$: a deferred MLP accounts for view-dependent effects. (3) $\psi^{k}$: a proposal MLP for sampling.
4.2 Consistent Training and Rendering
It is essential to ensure consistency between the training and rendering stages to achieve high-fidelity rendering results on the web as obtained during training. Due to the limited number of texture units within the web rendering environment, we are compelled to create multiple shaders to render distinct blocks. Specifically, one shader is allocated for storing the texture of an individual block. Each block subsequently renders an image respective to the current camera view. However, a simplistic averaging of these resultant rendering outputs can lead to discernible seams and does not ensure 3D consistency at the inter-block boundaries.
To address this problem, we simulate the process of multiple shaders rendering images and then linearly weighting them together using the volume rendering weights of blocks. For ray $\boldsymbol{r}(t)$, we uniformly sample between the near and far boundaries based on the scene’s bounding box. Then, according to the sample coordinates, we query the corresponding block’s proposal MLP to obtain the sample density, which is transformed into probability distributions along the rays. These probabilities guide a resampling strategy, ensuring a concentration on near-surface features with a few samples.
Figure 2: Visualization of alpha blending during consistent training. The left image shows the rendering results of four separate blocks by their shaders and the combined image after alpha blending, demonstrating how alpha blending correctly handles occlusion relationships. The right image visualizes the weights of rendering sample points before and after alpha blending. This illustration shows the correct occlusion management achieved through depth sorting of blocks followed by alpha blending.
@@ 4.2 Consistent Training and Rendering
Assuming that the proposal MLP yields samples passing through $\boldsymbol{M}$ blocks with a total of $\boldsymbol{N}$ samples, where each block $k$ has $\boldsymbol{n}_{k}$ samples, we simulate web rendering by performing volume rendering within each block to obtain its individual rendering diffuse color $\boldsymbol{C}_{d}^{k}$, specular feature $\boldsymbol{F}^{k}$ and opacity ${\alpha}^{k}$ according to  Eqs. 1 and 2. Then we get block $k$ final rendering color $\boldsymbol{C}^{k}$ according to Eq. 3. Consequently, for the sake of 3D consistency in rendering, we depth-sort the blocks and apply volume rendering across multiple blocks in sequence, using opacity to generate the volume rendering weights:
$\boldsymbol{C}(\boldsymbol{r})=\sum_{k}^{M}\prod^{k-1}_{j=1}(1-\alpha^{j})\boldsymbol{C}^{k}.$ (4)
@@
Under the Lambertian surface setting where the specular color is zero, the diffuse color and feature vector obtained from volume rendering on the total of $\boldsymbol{N}$ ray samples from Eq. 2 are equal to the results produced by our approach of conducting volume rendering within each block followed by inter-block volume rendering Eq. 4. The proof is given in the supplementary. Thus, our rendering approach maintains 3D consistency and simulates multiple shader rendering on the web as shown in Fig. 2.
During rendering, for the sample point $p_{i}$, we need to access the voxel grid corner and triplane grid corner where the point is located and use interpolation to obtain the sample point’s attributes. During training, we also simulate the voxel and triplane grid points to maintain consistency with the rendering process. By using grid corners’ positions to query the attribute query function $f_{k}$, we obtain the attributes of the grid corners. Through interpolation, we acquire the attributes of the sample points in a manner similar to the rendering pipeline. This simulaition strategy ensures that the values used for volume rendering within each block during training are as closely matched as possible to the values queried from the baked textures.

### v2

# https://ustc3dv.github.io/City-on-Web/

@@ Abstract
Existing neural radiance field-based methods can achieve real-time rendering of small scenes on the web platform. However, extending these methods to large-scale scenes still poses significant challenges due to limited resources in computation, memory, and bandwidth. In this paper, we propose City-on-Web, the first method for real-time rendering of large-scale scenes on the web. We propose a block-based volume rendering method to guarantee 3D consistency and correct occlusion between blocks, and introduce a Level-of-Detail strategy combined with dynamic loading/unloading of resources to significantly reduce memory demands. Our system achieves real-time rendering of large-scale scenes at approximately 32FPS with RTX 3060 GPU on the web and maintains rendering quality comparable to the current state-of-the-art novel view synthesis methods.
Keywords:
real-time rendering neural rendering large-scale reconstruction
@@ 1 Introduction
Neural radiance field (NeRF) has significantly advanced the field of scene reconstruction, showing an unparalleled ability to capture complex details across diverse environments. Existing works have demonstrated its ability to render small scenes with exceptional quality and performance in real-time [2, 46, 47, 37, 20, 29, 5, 40, 9, 43, 48, 24]. NeRF has also been successfully applied to the rendering of large scenes in offline settings, achieving exceptional visual fidelity and generating intricately detailed results [35, 11, 43, 45, 39].
Despite these successes, real-time neural rendering of large scenes on the web remains profoundly challenging due to inherent computational power, memory, and bandwidth limitations on commodity devices. MERF [30] has recently achieved significant progress by employing a baking technique to reduce query network calls in the rendering pipeline, thereby enabling real-time rendering of small-scale scenes on the web. However, MERF struggles to capture intricate details in large scenes due to its limited resolution. A naive solution would be to simply increase the volumetric representation’s resolution, but this approach would lead to unacceptable increases in memory usage, scaling with $O(N^{3})$, and a significant decrease in rendering speed.
To overcome these limitations, we integrate MERF with a block-based strategy [39] for reconstructing large scenes, a method supported by numerous studies [35, 39, 50]. This approach not only improves the model’s representational ability but also controls memory growth at an $O(N^{2})$ rate because we divide the scene based on ground coordinates without dividing the height. However, there are certain challenges associated with a resource-independent block-based rendering approach on the web. Specifically, rendering on the web faces limitations on the number and resolution of texture units that can be loaded into a shader (typically no more than 16 texture units), which prevents loading all block resources into a single shader. Consequently, we load the rendering resources of different blocks into their respective shaders. Nevertheless, rendering with different shaders causes issues with 3D consistency. Specifically, when a ray traverses multiple blocks, sampling points might belong to different blocks loaded by different shaders, preventing standard volume rendering. We are thus compelled to render each block sequentially and subsequently combine the rendering results of the different blocks. To this end, we propose a block-based volume rendering strategy and demonstrate that this method of sequential block rendering is equivalent to volume rendering, thereby ensuring correct occlusion and 3D consistency of the rendering results.
Moreover, when viewing from a higher altitude viewpoint, the rendering resources of all scene blocks are needed. Nonetheless, loading all blocks for rendering is impractical due to the excessive memory usage that would surpass the capacity of standard consumer devices. To address this issue, we draw inspiration from traditional graphics techniques [6, 7, 8, 12, 18, 23, 14] to create Level-of-Detail (LOD) for each block’s resources, dynamically selecting resources for rendering based on the camera’s position and field of view. This approach significantly reduces the resource demands during rendering, paving the way for smoother user experiences even on less capable devices.
In summary, the contributions of this paper include the following aspects:
  • $\bullet$
    We propose a block-based multi-shader volume rendering method, ensuring real-time high-fidelity rendering of large-scale scenes.
  • $\bullet$
    We employ an LOD strategy and dynamic loading/unloading strategies to adaptively manage rendering resources and significantly reducing the quantity of resources loaded and ensuring efficient resource utilization for large-scale scene rendering.
  • $\bullet$
    Our experiments demonstrate that our system achieves real-time rendering of large-scale scenes at approximately 32FPS with 1080P resolution on an RTX 3060 GPU, while maintaining rendering quality comparable to the current state-of-the-art (SOTA) methods for large-scale scenes.
2 Related Work
Figure 1: Overview of City-on-Web pipeline. During the training phase, we uniformly partition the scene and reconstruct it at the finest LOD. To ensure 3D consistency, we use a resource-independent block-based volume rendering strategy (Sec. 4.2). For LOD generation, we downsample virtual grid points and retrain a coarser model (Sec. 4.4). This approach supports subsequent real-time rendering by facilitating the dynamic loading of rendering resources.
Large-scale Scene Reconstruction. For radiance field reconstruction of large-scale scenes, a key issue lies in enhancing the model’s ability to adequately capture and render extensive scenes. Some works [35, 39, 10] address this by adopting a divide-and-conquer strategy, segmenting expansive scenes into smaller blocks, and applying localized NeRF processing to each. This approach significantly improves both the reconstruction quality and the model’s scalability to larger scenes. Switch-NeRF [50] employs a gating network to dispatch 3D points to different NeRF sub-networks. Grid-NeRF [45] utilizes a compact multiresolution feature plane and combines the strengths of smoothness from vanilla NeRF with the local detail capturing ability of feature grid-based methods [26, 4, 32], efficiently reconstructing large scenes with fine details. NeRF++ [48] enhances the reconstruction of unbounded scenes through its innovative multi-spherical representation. On the other hand, Mip-NeRF 360 [1] introduces a scene contraction function to effectively represent scenes that extend to infinity, addressing the challenge of vast spatial extents. F2-NeRF [41] takes this a step further by implementing a warping function for local spaces, ensuring a balance of computational resources and training data across different parts of the scene.
Real-time Rendering. Early works mainly focus on the real-time rendering of a simple single object. NSVF [20] improves NeRF by introducing a more efficient sparse voxel field, significantly accelerating rendering speed while maintaining high-quality output. KiloNeRF [29] utilizes thousands of small MLPs, each responsible for a tiny scene region, significantly reducing network evaluation time. In contrast, SNeRG [13] leverages pre-computed sparse grids, allowing for direct retrieval of radiance field information without needing network evaluation. Termi-NeRF [28] terminates ray marching in less impactful scene regions, slashing computation time. DONeRF [27] focuses on one sample using a depth oracle network, speeding up rendering while preserving scene quality. Recently, there have been developments that enable real-time rendering of neural radiance fields in small scenes. MERF [30] improves upon SNeRG by utilizing a voxel and triplane hybrid representation to reduce memory usage. MobileNeRF [5] introduces the polygon rasterization rendering pipeline, running NeRF-based novel view synthesis in real-time on mobile devices. BakedSDF [46] bakes volumetric representation into meshes and utilizes spherical harmonics for representing view-dependent color, while NeRF2Mesh [37] iteratively refine both the geometry and appearance of the mesh. Furthermore, several methods [38, 42] exploit the real-time rendering attributes of mesh representations alongside the robust representational potential of volume representations, particularly for rendering hair, translucent materials, and similar entities. These hybrid methods facilitate the achievement of high-fidelity real-time rendering. Recently, 3D Gaussian splatting [16] achieves real-time rendering by utilizing a novel 3D Gaussian scene representation and a rasterization-based rendering pipeline. However, extending this representation to large scenes is challenging due to its substantial memory consumption.
Level of Detail. Substantial works are devoted to integrating LOD methods into the fabric of traditional computer graphics [6, 7, 8, 12, 18, 23, 14, 22], aiming to streamline rendering processes, reduce memory footprint, bolster interactive responsiveness. Recently, some works have begun to apply LOD to neural implicit reconstruction. NGLoD [34] represents LOD through a sparse voxel octree, where each level of the octree corresponds to a different LOD, allowing for a finer discretization of the surface and more detailed reconstruction as the tree depth increases. Takikawa et al. [33] efficiently encode 3D signals into a compact, hierarchical representation using vector-quantized auto decoder method. BungeeNeRF [43] employs a hierarchical network structure, where the base network focuses on learning a coarse representation of the scene, and subsequent residual blocks are tasked with progressively refining this representation. TrimipRF [15] and LOD-Neus [51] leverage multi-scale triplane and voxel representations to capture scene details at different scales, effectively implementing anti-aliasing to enhance the rendering and reconstruction quality.
3 Background
Our exploration begins with an in-depth analysis of two influential works, SNeRG [13] and MERF [30], which have both set benchmarks for real-time rendering of the radiance field. SNeRG precomputes and stores a Neural Radiance Fields model in a sparse 3D voxel grid. Each active voxel in SNeRG contains several attributes: density, diffuse color, and specular feature vector that captures view-dependent effects. Additionally, an indirection grid is used to enhance rendering by either indicating empty macroblocks or pointing to detailed texels in a 3D texture atlas. This representation allows real-time rendering on standard laptop GPUs.
The indirection grid assists in raymarching through the sparse 3D grid by passing empty regions and selectively accessing non-zero densities $\sigma_{i}$, diffuse colors $\boldsymbol{c}_{i}$, and feature vectors $\boldsymbol{f}_{i}$ from baked textures. Integrating along each ray $\boldsymbol{r}(t)=\boldsymbol{o}+t\boldsymbol{d}$, they compute the sum of the weights, which can be considered as the pixel’s opacity:
The step size $\delta_{i}$ during ray marching is equal to the voxel width for an occupied voxel. The color $\boldsymbol{C}_{d}(\boldsymbol{r})$ and specular feature $\boldsymbol{F}_{s}(\boldsymbol{r})$ along the ray are accumulated using the same weights to compute the final diffuse color and specular feature of ray:
Subsequently, the accumulated diffuse color and specular feature vector, along with the positional encoding $PE(\cdot)$ of the ray’s view direction, are concatenated to pass through a lightweight deferred MLP $\Phi$ to produce a view-dependent residual color:
While SNeRG achieves impressive real-time rendering results, its voxel representation demands substantial memory, which poses limitations for further applications. MERF presents a significant reduction in memory requirements in comparison to extant radiance field methods like SNeRG. By leveraging hybrid low-resolution sparse voxel and 2D high-resolution triplanes, MERF optimizes the balance between performance and memory efficiency. Moreover, it incorporates two pivotal strategies to bridge the gap between training and rendering performance. MERF simulates finite grid approach during training, querying MLPs at virtual grid corners and simulates quantization during training to mimic the rendering pipeline closely.
4 Method
In this section, we present a method for representing and rendering large scenes on the web. Our approach utilizes a block and LOD strategy for rendering large-scale scenes, as described in (Sec. 4.1). A block-based volume rendering approach is introduced for the seamless blending of different blocks, utilized both during the training and rendering stages to ensure consistency (Sec. 4.2). Sec. 4.3 details the optimization strategies. The generation and refinement of LODs Sec. 4.4 are also explained.  Sec. 4.5 elucidates the baking strategy suitable for the representation.
4.1 Large-scale Radiance Field
Although real-time rendering methods like MERF can achieve high-quality real-time rendering for small-scale scenes, they face representational capacity challenges when applied to larger scenes. As mentioned in Sec. 1, utilizing a single MERF model to represent vast scenes is problematic due to its limited resolution, especially in terms of detailed and accurate reconstruction. Therefore, we represent scenes using multiple blocks. However, this approach necessitates employing an LOD strategy to reduce the number of resources that need to be loaded during the rendering phase. Thus, we adopt a block-based and LOD strategy for representing the whole scene in the rendering stage. We will elaborate on the representation used in the rendering stage and provide the representation used to reconstruct the scene in the finest LOD in the training stage.
Training Stage. In the training stage, we only represent and optimize the finest LOD. We uniformly partition the entire scene into $K$ blocks $\{\mathcal{B}_{k}\}_{k=1}^{K}$, each centered at $\mathbf{c}_{k}=(x_{k},y_{k})$, on the $xy$ plane (i.e., the ground plane). This approach stems from the observation that large scenes typically exhibit smaller scales in the $z$-direction compared to the $xy$-plane, prompting us to partition based on the ground plane and avoid subdividing along the $z$-axis. For a point $\mathbf{p}=(p_{x},p_{y},p_{z})\in\mathbb{R}^{3}$, we determine its corresponding block $\mathcal{B}_{k}$ based on its $xy$ coordinates, denoted as $\mathbf{p}_{proj}=(p_{x},p_{y})$:
$\mathbf{p}\in\mathcal{B}_{k},k=\mathop{\arg\min}\limits_{k}\|\mathbf{p}_{proj}-\mathbf{c}_{k}\|_{\infty}$ (4)
Within block $k$, the following trainable components are introduced: (1) $f^{k}$ is an attribute query function that adopts a hash encoding and an MLP decoder that outputs attributes of points such as densities, diffuse color and specular feature (2) $\Phi^{k}$ is a tiny deferred MLP account for view-dependent effects. (3) $\psi^{k}$ is a proposal MLP for sampling.
Rendering Stage. In the rendering stage, our scene representation includes hierarchical $L$ LODs representation for the scene. Specifically, as shown in the right figure, we merge $2\times 2$ blocks into one block between two consecutive LODs. As a result, for LOD $l$, where $l\in\{1,2,\ldots,L\}$, there are $K/4^{l-1}$ blocks. In each block, the following baked textures are used for rendering: (1) $f^{k,l}$ is an attribute query function that takes the coordinates of a sample point as input and directly accesses the opacity, diffuse color and specular features of the sample point from the baked sparse voxel and triplane textures. (2) $\Phi^{k,l}$ represents a tiny deferred MLP that accounts for view-dependent effects. (3) $\psi^{k,l}$ is used as a multi-level occupancy grid for sampling.
4.2 Block-based Volume Rendering
In the rendering stage, we create multiple shaders to render distinct blocks. Specifically, one shader is allocated for storing the texture of an individual block. Each block subsequently renders an image respective to the current camera view. However, a simplistic averaging of these resultant rendering outputs can lead to discernible seams and does not ensure correct occlusion at the inter-block boundaries as shown in Fig. 2(a). Therefore, we employ a block-based volume rendering strategy and combine it with depth sorting followed by alpha blending to ensure seamless boundaries and correct occlusion at the edges.
(a) Incorrect blending
@@ 4.2 Block-based Volume Rendering
4.6cm0.6pt1mm 1mm
(b) Ours alpha blending
Figure 2: Visualization comparison between the alpha blending method and others. (a) Top image: incorrect occlusion without depth sorting. Bottom image: incorrect rendering results when simply using $\alpha_{i}/(\sum_{j}\alpha_{j})$ as blending weights. (b) Left: rendering results of four separate blocks and the final blending result. Right: visualization of sample points’ rendering weights before and after alpha blending.
@@
Specifically, in the rendering stage, for a ray $\boldsymbol{r}(t)$ passing through $\boldsymbol{M}$ blocks with a total of $\boldsymbol{N}$ samples, where each block $k$ has $\boldsymbol{n}_{k}$ samples, we perform volume rendering within each block to obtain its individual rendering diffuse color $\boldsymbol{C}_{d}^{k}$, specular feature $\boldsymbol{F}^{k}$ and opacity ${\alpha}^{k}$ of the ray in block $\mathcal{B}_{k}$ according to  Eqs. 1 and 2. Then we get final rendering color $\boldsymbol{C}^{k}$ of block $\mathcal{B}_{k}$ according to Eq. 3. Subsequently, to correctly handle occlusion in rendering, we depth-sort the blocks and apply volume rendering across multiple blocks in sequence, using opacity to generate the blending weights:
$\boldsymbol{C}(\boldsymbol{r})=\sum_{k=1}^{M}\prod^{k-1}_{j=1}(1-\alpha^{j})\boldsymbol{C}^{k}.$ (5)
Figure 3: Block-based volume rendering. "DR" denotes deferred rendering. $\Phi$ represents the deferred MLP.
@@
Under the Lambertian surface setting where the specular color is zero, the color obtained from volume rendering on the total of $\boldsymbol{N}$ ray samples from Eq. 2 is equal to the results produced by our approach of conducting volume rendering within each block followed by inter-block volume rendering Eq. 5. The proof is given in the supplementary. Thus, our rendering approach maintains correct occlusion and keeps 3D consistency when using multiple shaders rendering on the web as shown in Fig. 2(b).
The volume rendering process in MERF involves integrating all sample points together, followed by deferred rendering. In contrast, as shown in Fig. 3, our block-based volume rendering is fundamentally based on segmented integration. Without the deferred rendering process, it is entirely equivalent to traditional volume rendering. However, if we adhere to the MERF rendering pipeline during the training process, it will lead to discrepancies between the rendering results during the training and rendering phases, ultimately affecting the rendering quality. To minimize this gap, we adopt the same rendering pipeline during the training stage as we do in the rendering stage.
Specifically, in the training stage, for ray $\boldsymbol{r}(t)$, we uniformly sample between the near and far boundaries based on the scene’s bounding box. Then, we determine that this point is inside $\mathcal{B}_{k}$ according to Eq. 4 and query the corresponding proposal MLP $\psi^{k}$ of $\mathcal{B}_{k}$ to sample probability distributions along the ray. Similarly, we also query the corresponding $f^{k}$ to obtain the attributes of the rendering sample points. Lastly, like in the rendering stage, we render each block sequentially to obtain the color and opacity for the ray in $\mathcal{B}_{k}$ and use  Eq. 5 to derive the final rendering result.